A marketing research firm wishes to compare the prices charged by two supermarket chains-Millers and Alberts. The research
Answer #1

Solution:

a. 9.73

b.2.11

c. Reject Ho

---------------------------------------------------------------------------------------------------------------------------

Explanation:-

Null Hypothesis (Ho): \mu_{1}=\mu_{2}

Alternative Hypothesis (Ha): 1 2

Test Statistics

(X1- X2) (1 2) t /n13/n2

(121.92 114.81) - (0) t /1.402/10 1.842/10

t = 9.725

Degrees of freedom, df is calculated using the formula

2 2 2 п, п, df 2 2 1 1 n1-1n n2-1n п,- п,

(1.42/10 1.842/10)2 df 1/9(1.42/10)21/9(1.842/10)2

df16.8 ~ 17

Using t-tables, the critical value at a/2 = 0.05/2 = 0.025 with 17 df is 2.11.

Since test statistics is greater than the critical value, we reject the null hypothesis.